OpenStudy (anonymous):

Name the property 1. (34 + 11) + 9 = (11 + 34) + 9 2. 25 + (–15 + 4) = (25 + (–15)) + 4 3. –(–21 + 8) = 21 + (–8) 4. –30 + (–16) = –16 + (–30)

OpenStudy (anonymous):

4. Associative

OpenStudy (anonymous):

I think 3 is Associative as well.

OpenStudy (anonymous):

@texaschic101

OpenStudy (abdullahm):

Associative property is a + (b + c) = (a + b) + c

OpenStudy (abdullahm):

Commutative property is a + b = b+ a

OpenStudy (abdullahm):

So @mlggamer do it again :)